# Loss Shaping Constraints for Long-Term Time Series Forecasting

### v1

@@ C.1 Extended experiment results
Figure 6: Full visualizations for the test loss landscapes of the Exchange, Weather and ECL datasets, on constrained and ERM losses.
C.1.1 Loss Shape Plots
Figure 6 shows the loss landscapes of the different transformer architectures tested on all three datasets. Some patterns in loss shapes across model architectures can be observed. For instance, Informer tends to have a jagged loss distribution with periodic discontinuities. On the other hand, Autoformer’s losses are concave or show seasonal patterns. These distributions could be due to the inherent inductive biases on each model architecture, such as the autocorrelation and trend decomposition mechanisms in Autoformer.
However, the particular shape of the loss remains idiosyncratic to each combination of dataset, model, and prediction window length. Furthermore, the effect of imposing constraints can be appreciated in multiple instances, where the constraints appear to have a corrective effect on extreme cases, improving MSE and MCV simultaneously.
Finally we also observe that, more often than not, loss landscapes are not linear, let alone constant. Moreover, in rare cases, such as Reformer on Exchange Rate with prediction window of 192 steps, the stepwise loss surprisingly trends downwards. This motivates further experimentation on different kinds of constraints that are better suited for a particular learning problem.
C.1.2 Correlation between training and test errors.
C.1.3 Correlations between prediction lengths

### v2

@@ D.2 Resilient Loss Shape Plots
D.4 Correlations Between Prediction Lengths
@@
D.5 Ablation Analysis for Constant Constraints
Model Train 25% Train 50% Train 75% Val 25% Val 50% Val 75%
Autoformer 48.85 39.92 45.48 17.64 46.98 2.61
FiLM 29.88 29.64 29.61 29.44 29.18 28.87
Informer 0.85 1.54 1.51 56.48 60.50 60.50
Nonstationary Transformer -35.28 -36.45 -39.60 -2.79 4.31 16.54
Pyraformer 7.82 14.63 14.63 15.89 21.22 23.40
Reformer 1.76 1.53 1.72 -41.37 -48.63 -50.12
Transformer -0.73 0.24 0.20 -2.46 53.18 54.38
iTransformer -11.75 -13.02 -13.91 -15.52 -8.76 -8.74
Table 8: MSE % changes of constrained runs relative to ERM for varying values of $\epsilon$, in the Exchange Rate dataset, with window length 720.
By examining the loss landscape across varying values of $\epsilon$ used on the Exchange Rate dataset, we find that our constrained method is not overly sensitive to the choice of hyperparameters. In Table 8, we can observe each model’s relative changes in MSE. Pyraformer, Reformer, Informer and Nonstationary Transformer exhibit a biphasic behavior: the loss shape changes drastically only after the $\epsilon$ is set to a parameter much higher than ERM’s average training loss, which results in a trivial constraint. In the cases of FiLM and iTransformer, the shape remains relatively consistent across all values, while Autoformer exhibits larger variations. The Resilient approach is less sensitive to the constraint level, in some cases allowing for better generalization with the same value of epsilon compared to the constrained run.
Figure 14 illustrates the effect of the $\epsilon$ parameter on various models on the Exchange Rate dataset. We choose $\epsilon$ to be the quartiles of the train and test error taken from an ERM run. Notably, for several models, including the Informer, Reformer, and Transformer, the loss landscape exhibits relative invariance across similar $\epsilon$ values. This suggests a plateau in sensitivity within this parameter range. Conversely, in some cases, a critical $\epsilon$ threshold seems to exist, after which the shape significantly changes and then stabilizes again as we relax $\epsilon$ further. Specifically, for Autoformer, Informer, and Transformer, optimal performance is achieved subsequent to a degree of $\epsilon$, underscoring the benefit of exploring multiple constraint levels.
We have shown that our method is reasonably robust to the choice of the constraint hyperparameter. However, we note that performing a parameter search with different values of $\epsilon$ may still be valuable. When the generalization gap is high, training with constraints higher than all of the training values is often more effective. For example, Reformer’s test MSE is 2.9x higher than its train MSE, and Table 8 shows its MSE can be reduced by up to 50.1% when using validation-based threshold values, while training-based ones are not as effective.
inline][JP]: Check if Reformer’s generalization gap in exchange is higher than other models inline][IH]: I think that the meaning of train based and val based is not clear here.
Figure 14: Ablation plots for the $\epsilon$ parameter on the Exchange Rate dataset, with predictive window length of 720. Rows are different models, and columns are increasing values of $\epsilon$. Each plot shows test MSE of Constrained and Resilient runs.
D.6 Outlier in STD histogram
In Section 5.1, we showed that on average, using constraints reduces Window STD. However, we identified some outliers in Figure 4. In one instance, the constrained Window STD is 371% higher than ERM. It corresponds to Reformer on the Weather dataset with predictive window length of 720 steps, shown in Figure 9. Closer inspection reveals that the reason for this increase is due to the constrained model finding a much lower MSE (constrained 0.788 versus ERM 0.963) in the first part of the window. Interestingly, the resilient model settles for a middle point, with a lower MSE than ERM but less variance increase than the constrained model.